how do i solve:
if y=2 2/3 when x=1/4 find y when x=1 1/8
im not sure what category direct variation is under but some help would be great thank you :)

Answer by Cromlix(4381)

Hi there,
If y = 2 2/3 when x = 1/4
2 2/3 = 8/3
x = 1/4 when y = 8/3
Divide both by 1/4
x = 1 when y = 32/3
When x = 1 1/8 = 9/8
y = 32/3 x 9/8
y = 12.
Hope this helps :-)